# A Survey on Speech Large Language Models

### v2

@@ Abstract
Large Language Models (LLMs) exhibit strong contextual understanding and remarkable multi-task performance. Therefore, researchers have been seeking to integrate LLMs in the broad sense of Spoken Language Understanding (SLU) field. Different from the traditional method of cascading LLMs to process text generated by Automatic Speech Recognition(ASR), new efforts have focused on designing architectures centered around Audio Feature Extraction - Multimodal Information Fusion - LLM Inference (Speech LLMs). This approach enables richer audio feature extraction while simultaneously facilitating end-to-end fusion of audio and text modalities, thereby achieving deeper understanding and reasoning from audio data. This paper elucidates the development of Speech LLMs, offering an in-depth analysis of system architectures and training strategies. Through extensive research and a series of targeted experiments, the paper assesses Speech LLMs’ advancements in Rich Audio Transcription and its potential for Cross-task Integration within the SLU field. Additionally, it indicates key challenges uncovered through experimentation, such as the Dormancy of LLMs under certain conditions. The paper further delves into the training strategies for Speech LLMs, proposing potential solutions based on these findings, and offering valuable insights and references for future research in this domain, as well as LLM applications in multimodal contexts.
@@ 1 Introduction
The advancements in large language models within natural language processing have been particularly significant, especially in contextual understanding and multimodal technology. Models like GPT-4 [1], trained on extensive corpora, are now capable of processing complex contexts with exceptional accuracy. These models, with their vast parameters, excel in parsing contextually appropriate sentences and maintaining coherence over multiple conversational turns, which is crucial for tasks like dialogue systems, automatic summarization, and machine translation. By leveraging vast contextual information, these models manage long-text processing with unprecedented proficiency in contextual reasoning.
In addition to the success in text-based tasks, large language models have also achieved remarkable success in multimodal tasks [2], particularly in integrating and processing inputs from diverse modalities like vision, speech, and text. Their ability to semantically link these different types of data has led to significant breakthroughs in areas such as visual question answering and image generation, as exemplified by models like CLIP [3] and DALL-E [4]. This capability not only enhances performance in visual tasks by effectively fusing information from different modalities, but also demonstrates the models’ potential to expand into and improve upon other modalities. The way these models manage complex multimodal information processing highlights their adaptability and lays the groundwork for further innovations in multimodal understanding and integration.
The typical Spoken Language Understanding (SLU) involves interpreting spoken language to extract meaning, intent, and relevant information beyond simple transcription [5, 6]. In order to more comprehensively and coherently demonstrate the performance of Speech LLMs, in this article, we also include basic speech-to-text tasks, such as automatic speech recognition (ASR), within its scope. Related works have achieved remarkable performance across various domains, including voice-activated assistants, automated transcription, and real-time speech translation. Modern systems are adept at handling diverse accents and languages, improving efficiency and accuracy in workflows like medical transcription and customer service automation. However, challenges remain, particularly in long-form speech recognition and hotword/keyword recognition. Long-form recognition struggles with maintaining context over extended periods and handling speaker variability, often resulting in accuracy degradation and latency issues in real-time applications. Hotword detection, critical for wake word recognition, faces difficulties in noisy environments and balancing sensitivity and specificity, especially when hotwords are contextually similar to other phrases. The integration of hotword recognition into continuous speech processing without disrupting flow or increasing computational overhead remains an ongoing challenge.
In this paper, our contributions can be summarized in the following three points:
@@
    The first in-depth and comprehensive survey analyzing Speech LLMs in the SLU domain. We will conduct an in-depth exploration of the development of Speech LLMs, explaining the paradigm of model architecture, and perform a comparative analysis with traditional speech models in SLU tasks.
@@
    The first comparative analysis of training methods for aligning speech and text modalities in Speech LLMs architecture, with an emphasis on the potential development of Reinforcement Learning strategies such as DPO and PPO.
@@ 2.1 Evolution of Speech LLMs Architectures
Considering the exceptional performance of large language models (LLMs) in contextual understanding, reasoning, and multimodal information processing, and recognizing the unresolved challenges in the field of SLU, such as long-form speech understanding and hotword recognition, researchers have begun to explore the integration of LLMs into SLU.
@@ 2.1.1 Transformer Integration into Traditional Speech Models
One of the key milestones in the integration of LLMs into SLU was the introduction of the Transformer architecture. Dong et al. (2018) first applied the Transformer to speech recognition, proposing an end-to-end speech recognition system based on the Transformer [7]. This attempt laid the foundation for subsequent research. Following this, in 2020, Gulati et al. (2020) further advanced this approach by introducing the Conformer architecture, which combines the local feature extraction capabilities of Convolutional Neural Networks (CNNs) with the Transformer, making the Transformer more robust in processing speech signals [8].
In 2021, Hsu et al. proposed HuBERT (Hidden-Unit BERT, which represented a major breakthrough in utilizing LLMs to process audio features [9]. HuBERT employs self-supervised learning on a large corpus of unlabeled speech data to learn effective hidden unit representations. In 2022, Radford et al. introduced the Whisper model, which further integrated multilingual and multitask capabilities into a single model, demonstrating excellent transferability across different languages and tasks [3].
Researchers, unsatisfied with simply using speech as input and text as output, began exploring end-to-end speech-text multimodal inputs and outputs. In 2022, Junyi Ao et al. introduced SpeechT5, pioneering the investigation of a unified encoder-decoder framework for various spoken language processing tasks [10].
@@ 2.1.2 Direct Audio Processing with Large Language Models
Fig. 1: The evolution of Speech LLMs from the Pre-LLM Stage with separate acoustic and language models, to the Early LLM Stage with transformer integration, and the Current Speech LLM Stage, featuring direct audio processing by large language models. The current stage includes two branches: Discrete Sequence Modeling (e.g., VALL-E, SpeechGPT) and Continuous Sequence Modeling (e.g., Pengi, SALMONN).
@@
Besides directly incorporating the architecture of LLMs like transformer into speech recognition tasks, another significant integration approach involves using LLM as a whole to process audio features extracted by traditional speech recognition tools. This approach leverages the powerful contextual understanding and reasoning abilities of LLMs to improve the accuracy and robustness of speech recognition and deeper speech understanding, and it has gradually evolved into the main trend of using large language models in the field of speech recognition.In other words, this refers to the narrower definition of Speech LLMs .
The earliest attempts to align speech and text modalities by passing extracted features from speech to downstream language models can be traced back to 2020. Qi Liu et al. were the first to explore this approach, separating the mapping of speech feature information from the speech feature extraction process into two independent modules and introducing the LSTM model into the process of mapping speech feature information to text [11]. This approach differs from traditional CTC-based end-to-end systems by isolating the acoustic feature-to-text mapping as a standalone unit. This separation laid the groundwork for the subsequent integration of large language models into this area, making it possible to embed LLMs more effectively in speech recognition tasks.
Currently, Speech LLMs can mainly be divided into two categories 4. The first is Discrete Sequence Modeling, represented by approaches that condense audio feature information into discrete tokens, which are then passed to the LLM for processing. In early 2023, Microsoft released Vall-E, a typical study utilizing LLMs to process audio features. Vall-E introduces the Transformer architecture, combining audio features with the capabilities of language models to achieve more natural text-to-speech generation [12]. Following this, the emergence of SpeechGPT marked a deep integration of speech models and LLMs, capable not only of processing audio but also interacting through natural language, bringing a new interactive paradigm to the field of speech recognition [13]. In mid-2023, Google launched AudioPaLM, which further expanded the capabilities of speech recognition into multimodal processing, enhancing performance in multimodal tasks [14].
The second is Continuous Sequence Modeling, where audio feature information is projected into continuous input embedding vectors and then transmitted to the LLM for processing. In 2023, Deshmukh et al. introduced Pengi, which projects speech modality information into the text space of large models without altering any parameters in LLM such as embedding vector spaces [15]. Subsequently, continuous audio feature information began to be used to convey richer representations to the LLM [16, 17, 18], resulting in significant performance improvements across a range of tasks in the SLU domain. In 2024, Ma et al. further proposed SLAM-LLM, which involves the addition of a linear projector, allowing the task to be greatly completed by training only the projection layer [19].
Once researchers achieved significant improvements in recognition accuracy compared to traditional speech recognition technologies, they began exploring and attempting to address the issues inherent in these conventional methods. In 2024, Fathullah et al. introduced the Conformer mechanism, achieving advancements in handling long speech sequences [20]. In the same year, SpeechX achieved breakthroughs in multilingual, multitask speech recognition, enabling seamless switching between languages and supporting challenges like long-form speech understanding and hotword recognition [21].
2.2 Advancements of Speech LLMs in Key Tasks and Challenges
With the continued advancement and application of large language models (LLMs) in the field of speech processing, the Speech LLM paradigm has demonstrated significant success across various tasks, often surpassing the performance of traditional models in many areas. Our primary focus is on traditional key tasks within the domain of speech understanding, which include automatic speech recognition (ASR), speech translation, and speaker identification, among others. A more detailed comparison and analysis will be presented in  Section 6. At the same time, we also focus on the exploration and breakthroughs of Speech LLM in challenging tasks such as long-form speech recognition, as well as keyword and hotword detection. Of course, the powerful contextual reasoning capabilities of large language models provide an alternative solution for more complex system tasks. In this regard, we have also explored areas such as real-time multimodal interaction. This section investigates both the performance of Speech LLMs on these standard tasks and how they address long-standing challenges in the field.
2.2.1 Improvement in Traditional Tasks in Spoken Language Understanding
Traditional tasks in the field of speech understanding include automatic speech recognition, speaker identification, and speech translation. The Speech LLM paradigm has demonstrated impressive performance in all these areas. Here, we will firstly take ASR as an example to briefly examine and understand the progress Speech LLM has made in this field.
Automatic Speech Recognition (ASR) is the technology that enables machines to convert spoken language into text. Modern ASR systems, enhanced by LLMs, aim to achieve higher accuracy, better noise resilience, and greater adaptability to diverse accents and dialects compared to their predecessors. This task is foundational for voice-controlled applications, interactive voice response systems, and automated transcription services. One of the key metrics for evaluating speech recognition systems is the Word Error Rate (WER).
Traditional models, such as those based on LSTM or GRU, have shown competitive performance on various benchmark datasets [22]. However, the introduction of LLMs has significantly improved these results [23, 24].
Moreover, in multilingual speech recognition, LLMs have demonstrated superior performance across various languages. On the Multilingual LibriSpeech (MLS) dataset, Conformer-based models have achieved significant reductions in WER across multiple languages [25]. Notably, in low-resource languages such as French and German, great WER reductions have been reported [23].
Speech translation is the process of converting spoken language from one language into written or spoken text in another language. It typically involves two key steps [26]: automatic speech recognition (ASR), which transcribes spoken words into text, and machine translation (MT), which translates the transcribed text into the target language. This technology is used in real-time applications such as multilingual meetings, conferences, and live broadcasts, enabling seamless cross-lingual communication. With the successful application of large language models in the field of machine translation [27], the speech translation domain has also begun to gradually incorporate large language models, aiming to achieve similar significant results. Advances in large language models have significantly improved the accuracy and fluency of speech translation systems. In the field of speech translation, the integration of large language models has significantly enhanced performance. Not only has it improved the accuracy of speech translation tasks [28], but it has also broadened the range of supported languages [29], making these models an invaluable addition to the field. The incorporation of large language models brings a complementary advantage, enriching both precision and diversity in multilingual speech translation.
Additionally, LLMs have excelled in multitask learning scenarios. For example, the Qwen-Audio model has shown impressive performance in tasks that combine speech-to-text with other modalities, such as sentiment analysis and speaker identification, reducing WER by 10% and improving sentiment recognition accuracy by 8% compared to single-task models [17].
For a detailed comparison and analysis of Speech LLMs’ performance in traditional speech tasks, please refer to Section 6.
2.2.2 Long-Form Speech Understanding
Traditional speech recognition systems have struggled with long-form speech understanding due to context loss over extended periods. This issue is particularly pronounced in audio segments longer than one minute, where traditional models often exhibit a sharp increase in WER. In contrast, LLM-based systems have significantly mitigated this problem. The Whisper model, for instance, maintains contextual consistency across long-form audio, demonstrating an 18% reduction in WER on audio segments exceeding five minutes compared to traditional models. This enhancement is critical for applications such as the transcription of speeches and lectures.
Additionally, models like UniAudio and Pengi have shown remarkable performance in maintaining low WERs across extended speech segments by integrating advanced contextual understanding [30, 31]. These models have been particularly effective in domains requiring detailed, accurate transcriptions of prolonged speech events, where maintaining context over extended durations is crucial.
2.2.3 Hotword Recognition
Hotword recognition has been another challenging area for traditional speech recognition systems, especially in noisy environments. Traditional models often fail to accurately detect and differentiate hotwords within long audio streams. The GenTranslate model, introduced in late 2023, has made substantial improvements in this area by leveraging the contextual understanding capabilities of LLMs [28]. In a test set containing multiple domain-specific hotwords, GenTranslate achieved a 22% improvement in hotword recognition accuracy compared to traditional models. Moreover, it maintained high robustness in noisy conditions, marking a significant advancement over prior approaches.
Further examples of LLM-based models outperforming traditional systems include Mala-ASR and Whisper, which not only improves hotword recognition accuracy [32] but also adapts dynamically to new hotwords in real-time, a feature that traditional models lack. This capability is particularly valuable in dynamic environments like live broadcasts or interactive voice response (IVR) systems where new terms may frequently be introduced.
2.2.4 Real-time Multimodal Interaction
The integration of LLMs into speech recognition has expanded the scope of tasks beyond traditional speech-to-text, enabling real-time multimodal interaction. Models like VoxtLM and LauraGPT facilitate seamless integration of speech with visual and textual inputs [33, 34], providing coherent and accurate multimodal outputs. This capability is particularly valuable in applications such as live transcription and synchronized translation during presentations, where both speech and visual context need to be processed simultaneously.
Moreover, LLM-based systems have introduced new functionalities, such as the generation of descriptive text, summaries, and even translations based on audio input. For instance, ViOLA can generate coherent summaries and cross-language translations with high fluency and accuracy, outperforming traditional models in both speed and quality of output. This represents a significant advancement in how speech recognition systems can interact with and interpret complex multimodal data streams [35].In addition, the integration of large language models with acoustics has demonstrated significant development potential across various specialized fields [36].
@@ 3 Model Paradigm of Speech LLMs
As a model architecture that integrates both audio and text modalities, Speech LLMs exhibits significant differences in input-output processing and model structure compared to traditional speech models. Notably, due to the significant commonality and prevailing trends, we choose to focus on the narrower definition of Speech LLMs 1, which involves the integration of LLMs as a whole into SLU tasks.
3.1 Speech LLM Input-Output Modes
Fig. 2: Overview of Speech LLM architectures corresponding to three different input-output modalities: Speech-to-Text (S2T), Speech&Text-to-Text (ST2T), and Speech&Text-to-Speech&Text (ST2ST).
Generally, the input and output modes of Speech LLMs can be categorized into three main types:

### v4

@@ Abstract
Large Language Models (LLM) exhibit strong contextual understanding and remarkable multitask performance. Therefore, researchers have been actively exploring the integration of large language models (LLM) into the domain of speech understanding, with a primary focus on a broad range of speech-to-text tasks. These include automatic speech recognition (ASR), speech-to-text translation (ST), speech emotion recognition (SER), and others. We refer to such models as Speech LLMs, which are typically built on a unified architecture following the pipeline of Audio Feature Extraction $\rightarrow$ Multimodal Information Fusion $\rightarrow$ LLM Inference. This approach enables richer audio feature extraction while facilitating end-to-end fusion of audio and text modalities, thereby achieving deeper understanding and reasoning from audio data. This paper elucidates the development of Speech LLMs, offering an in-depth analysis of system architectures. Through extensive research and a series of targeted experiments, the paper assesses the advancements in Speech LLMs and its potential for cross-task integration within the speech understanding field. Furthermore, it indicates key challenges discovered through experimentation, such as the dormancy of LLM under certain conditions. The paper further delves into the training strategies for Speech LLMs, proposing potential solutions based on these findings, and offering valuable insights and references for future research.
@@ 1 Introduction
The advancements in large language models within natural language processing have been particularly significant, especially in contextual understanding and multimodal technology. Models like GPT-4 [1], trained in extensive corpora, are now capable of processing complex contexts with exceptional accuracy. In addition to success in text-based tasks, large language models have also achieved remarkable success in multimodal tasks [2], particularly in integrating and processing input from various modalities such as vision, speech, and text. Their ability to semantically link these different types of data has led to significant breakthroughs in areas such as visual question answering and image generation, as exemplified by models like CLIP [3] and DALL-E [4].
In the field of speech processing, large language models have gradually begun to be widely applied [5, 6]. Instead of using Automatic Speech Recognition (ASR) technology to recognize cascaded large language models  [7, 8], the Audio Feature Extraction - Multimodal Information Fusion - LLM Inference has become the most widely adopted approach. Unlike other survey papers that might also focus on generative tasks in speech  [9], our work does not center on speech generation, but rather focuses primarily on their development and application in speech understanding tasks, including traditional tasks such as acoustic speech recognition and speech-to-text translation, as well as tasks such as speaker recognition, speech emotion recognition, and others. Speech LLMs have not only achieved significant success in the tasks mentioned above  [10, 11, 12], but also demonstrated impressive capabilities and potential in challenging tasks such as long-form speech recognition, hotword and keyword recognition, etc. [13, 14]
In this paper, our contributions can be summarized in the following three points.
@@
    The first in-depth and comprehensive survey that analyzes Speech LLMs. We will conduct an in-depth exploration of the development of Speech LLMs, explaining the paradigm of model architecture, and perform a comparative analysis with traditional speech models in speech understanding tasks.
@@
    The first comparative analysis of training methods for aligning speech and text modalities in Speech LLMs architecture, as well as the first to provide an in-depth classification and summary of the training datasets for Speech LLMs.
@@ 2 Recent Advances in Speech LLMs
2.1 Advancements of Speech LLMs in Speech Understanding
With the increasing integration of large language models into the speech understanding, the Speech LLMs paradigm has shown substantial progress across a range of speech understanding tasks. Rather than replacing existing solutions, Speech LLMs enhance them by combining rich audio representations with strong language reasoning capabilities.
In traditional tasks such as automatic speech recognition (ASR) and speech-to-text translation(ST), Speech LLMs consistently achieve higher accuracy than conventional models [15, 16]. The use of continuous representations enables these models to retain more fine-grained acoustic and contextual information, which is particularly beneficial in high-precision applications. Furthermore, Speech LLMs have demonstrated notable improvements in multilingual and low-resource scenarios, expanding the accessibility and robustness of speech technologies [17, 18].
Beyond accuracy, another key advantage of Speech LLMs lies in their multitasking capability. A single model can seamlessly perform various speech understanding tasks, including ASR, translation, speaker identification, emotion recognition, voice activity recogintion and even keyword spotting  [19, 20, 21, 12]. This unified modeling approach reduces system complexity and facilitates the development of more versatile speech systems.
Moreover, Speech LLMs might address long-standing challenges in the field. In long-form speech recognition, they help maintain contextual consistency over extended durations, significantly reducing error rates in processing lengthy audio. In keyword and hotword recognition, they offer higher precision and better robustness in noisy or dynamic environments by leveraging global context and adaptive modeling  [14, 22, 23].
Together, these advancements position Speech LLMs as a powerful foundation for building next-generation speech systems that are not only accurate but also adaptable and capable of handling diverse real-world tasks7.
@@ 2.2 Evolution of Speech LLMs Architectures
We broadly categorize the development of Speech LLMs into two phases. The first phase corresponds to the precursors of Speech LLMs, where the transformer architecture, having demonstrated significant success in natural language processing (NLP), was subsequently applied to traditional speech recognition and related models. The second phase involves integrating large language models (LLM) as a unified module following the speech feature extractor, establishing them as a critical component of speech processing.
@@ 2.2.1 Transformer Integration into Traditional Speech Models
Early explorations of integrating Transformers into speech models retained the traditional speech modeling paradigm, incorporating the Transformer architecture as a component rather than fundamentally altering the overall model structure. Dong et al. (2018) applied the Transformer to speech recognition, proposing an end-to-end speech recognition system based on the Transformer [24]. Following this, in 2020, Gulati et al. (2020) further advanced this approach by introducing the Conformer architecture, which combines the local feature extraction capabilities of Convolutional Neural Networks (CNNs) with the Transformer, making the Transformer more robust in processing speech signals [25].
In 2021, Hsu et al. proposed HuBERT (Hidden-Unit BERT) [26]. HuBERT employs self-supervised learning on a large corpus of unlabeled speech data to learn effective hidden unit representations. The encoder part of the transformer processes the audio features and generates contextualized representations, which are then used for the prediction task. In 2022, Radford et al. introduced the Whisper model, which further integrated multilingual and multitask capabilities into a single model. Whisper also leverages the power of the transformer architecture to handle the complex mappings between speech and text. The self-attention and cross-attention mechanisms of the transformer allow Whisper to attend to different parts of the speech signal when generating the text output, improving its accuracy and robustness [3].
@@ 2.2.2 Direct Audio Processing with Large Language Models
Fig. 1: The evolution of Speech LLMs from the Pre-LLM Stage with separate acoustic and language models, to the Early LLM Stage with transformer integration, and the Current Speech LLMs Stage, featuring direct audio processing by large language models. The current stage includes two branches: Discrete Sequence Modeling (e.g., VALL-E, SpeechGPT) and Continuous Sequence Modeling (e.g., Pengi, SALMONN).
@@
Besides directly incorporating the architecture of LLM like the transformer into speech recognition tasks, another significant integration approach involves using LLM as a whole to process audio features extracted by the audio feature extractor. This approach leverages the powerful contextual understanding and reasoning abilities of LLM to improve the accuracy and robustness of speech recognition and achieve a deeper understanding of speech, and it has gradually evolved into the main trend of using large language models in the field of speech understanding. In other words, this refers to the narrower definition of Speech LLMs .
Over time, researchers have converged on two primary strategies for integrating speech with large language models (Figure 4). The first, known as discrete sequence modeling, involves converting raw audio into a sequence of discrete units. These units, which capture essential features of the speech signal, allow the LLM to operate in a familiar token-based paradigm, bridging the gap between speech and text processing. Importantly, these discrete tokens are not only suitable for language modeling but also enable high-quality reconstruction of the original audio through neural vocoders. In early 2023, Microsoft released Vall-E, a zero-shot text-to-speech model exemplifying how EnCodec’s tokens are used in a Speech LLMs. By modeling acoustic tokens instead of waveform samples directly, VALL-E can generate high-quality speech that preserves the speaker’s identity, emotion, and acoustic environment from the prompt  [27]. Following this, the emergence of SpeechGPT marked a deep integration of speech models and LLM, capable not only of processing audio but also of interacting through natural language. By training on paired unit sequences and transcripts, the LLM learns to treat discrete HuBERT tokens as another language it can understand and generate. This allows SpeechGPT inherently “speak” and “listen”: the LLM can take spoken instructions that are converted to units and generate either a text or a sequence of units as a response [28]. In mid-2023, Google launched AudioPaLM, which further expanded the capabilities of speech recognition into multimodal processing by leveraging another major line of work in speech tokenization from the wav2vec family, improving performance in multimodal tasks [29, 30].
The second paradigm is Continuous Sequence Modeling, where audio features are projected into continuous embedding vectors that are directly consumed by the LLM. Compared to the discrete formulation, this approach offers a more direct and arguably more natural way to model speech, as it preserves the continuous nature of audio signals throughout the processing pipeline. In 2023, Deshmukh et al. introduced Pengi, which projects information about speech modality into the text space of large models without altering any parameters in LLM such as embedding vector spaces [31]. By preserving more of the rich information inherent in the audio signal, continuous embeddings enable Speech LLMs to achieve state-of-the-art performance on high-precision tasks such as ASR and ST [10, 32]. This paradigm’s ability to retain fine-grained acoustic details makes it particularly well-suited for applications where accuracy is critical. Moreover, these models also exhibit strong multitask capabilities, enabling them to generalize across various speech-related tasks [20, 19, 21]. This approach might require only the insertion of a simple projection layer between the audio encoder and the LLM and a minimal fine-tuning cost, yet it can achieve state-of-the-art performance on certain tasks [33, 15].
Researchers have not only pursued multitask capabilities and higher recognition accuracy beyond what traditional speech recognition technologies offer, but have also simultaneously sought to address long-standing challenges inherent in conventional speech understanding frameworks using this architecture [34, 14].
Considering the focus on the field of speech understanding, the second approach is more widely adopted. Therefore, this paper will primarily elaborate on the second approach while providing a comparative analysis with the first approach.
@@ 3 Model Paradigm of Speech LLMs
As a model architecture that integrates both audio and text modalities, Speech LLMs exhibits significant differences in input-output processing and model structure compared to traditional speech models. Notably, due to the significant commonality and prevailing trends, we choose to focus on the narrower definition of Speech LLMs 1, which involves the integration of LLM as a whole into Speech Understanding tasks.
3.1 Speech LLMs Input-Output Modes
Fig. 2: Overview of Speech LLMs architectures corresponding to three different input-output modalities: Speech-to-Text (S2T), Speech&Text-to-Text (ST2T), and Speech&Text-to-Speech&Text (ST2ST).
Speech LLMs can generally be categorized into three main types based on their input and output modes, with a particular focus on speech understanding tasks in this paper.: